# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v5

@@ 5 Efficient Algorithm for Computing the Optimal Bounds
In this section, we present an efficient algorithm to compute the optimal privacy amplification bounds within the general clone framework. Our approach builds upon the previously overlooked concept of the privacy amplification random variable (PARV), originally proposed in the privacy blanket framework [BBGN19]. While PARV offers an exact expression for the amplification bound, the original work did not provide a method for computing it precisely, relying instead on loose analytical approximations.
@@ 5.1 Generalized privacy amplification random variable
The following theorem is a remarkable result from the privacy blanket paper [BBGN19]. We restate it here in a more general form.
@@
We provide the proof in Appendix C. It is straightforward to observe that Theorem 12 can be directly applied to compute the general clone bound $D_{e^{\varepsilon}}(P_{0}^{\mathsf{GC}}\parallel P_{1}^{\mathsf{GC}})$. Consider the general clone induced by the primary optimal decomposition; the corresponding random variable $G$ can be expressed as follows:
@@
$\text{G}_{\varepsilon}^{x,x^{\prime}}$ =$=\begin{cases}\frac{1}{\gamma}L_{\varepsilon}^{x,x^{\prime}},&\text{w.p.}\ \gamma,\\ 0,&\text{w.p.}\ 1-\gamma.\end{cases}$
The quantity $L_{\varepsilon}^{x,x^{\prime}}$ was introduced by Balle et al. [BBGN19] and referred to as the Privacy Amplification Random Variable (PARV). We refer to $G_{\varepsilon}^{x,x^{\prime}}$ as the Generalized Privacy Amplification Random Variable (GPARV). The following theorem follows directly from Theorem 12 and the definition of the primary optimal decomposition in Section 4.2.
@@
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))\leq D_{e^{\varepsilon}}(P_{0}^{B}\parallel P_{1}^{B})=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
We now briefly revisit how the original privacy blanket paper utilizes the PARV. Using PARV, Balle et al. derived the precise expression for $D_{e^{\varepsilon}}(P_{0}^{B}\parallel P_{1}^{B})$:
Lemma 3 (Lemma 5.3 in [BBGN19]).
@@
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))\leq D_{e^{\varepsilon}}(P_{0}^{B}\parallel P_{1}^{B})$ $=\frac{1}{\gamma n}\sum_{m=1}^{n}\binom{n}{m}\gamma^{m}(1-\gamma)^{n-m}\mathbb{E}\left[\sum_{i=1}^{m}L_{i}\right]_{+}$
@@
In the original paper of the privacy blanket framework, a similar property of PARV was provided, but in a loose form [BBGN19]. Specifically, they established that $\gamma(e^{-\varepsilon_{0}}-e^{\varepsilon+\varepsilon_{0}})\leq L\leq\gamma(e^{\varepsilon_{0}}-e^{\varepsilon-\varepsilon_{0}})$. However, the $\varepsilon_{0}$-DP property of $\mathcal{R}$ actually guarantees a tighter bound: $\gamma(1-e^{\varepsilon+\varepsilon_{0}})\leq L\leq\gamma(e^{\varepsilon_{0}}-e^{\varepsilon})$.
@@ 5.2 Algorithm for computing privacy amplification upper bounds
The distributions of $G_{\varepsilon}$ for commonly used local randomizers are summarized in Table 1. Since the optimal decompositions and corresponding GPARVs of BLH, RAPPOR, and OUE depend on the domain size $D$ (see Table 2 in Appendix A), we present their GPARVs in the asymptotic regime (i.e., for $D\gg 1$) for simplicity. The distribution of $G_{\varepsilon}$ for Laplace mechanism on $\{0,1\}$ is presented in Appendix B. (The Laplace mechanism over the interval $[0,1]$ does not satisfy input symmetry.)
@@
Because $G$ is supported on the interval $[1-e^{\varepsilon_{0}+\varepsilon},e^{\varepsilon_{0}}-e^{\varepsilon}]$ (Property 2), the discretization step requires $O\left(\frac{e^{\varepsilon_{0}+\varepsilon}}{l}\right)$ operations. For most frequency oracles used on categorical data, $G$ takes values on at most five points (see Table 1), resulting in $O(1)$ discretization time.
@@
The above analysis concerns the error introduced when using the FFT-based algorithm to compute $D_{e^{\varepsilon}}\big{(}P_{0}^{B}\parallel P_{1}^{B}\big{)}$. From a theoretical perspective, since $\delta$ is typically very small, achieving high-precision computation (i.e., with error $o(\delta)$) would require a very small discretization step size $l$, which incurs a high computational cost.
However, our experimental results show that when discretization is used to infer the smallest value of $\varepsilon$ such that $D_{e^{\varepsilon}}\big{(}P_{0}^{B}\parallel P_{1}^{B}\big{)}\leq\delta$, the resulting error in $\varepsilon$ is negligible in practice. This is because the distribution of the GPARV is highly sensitive to changes in $\varepsilon$. In particular, once $\varepsilon$ exceeds the true threshold $\varepsilon^{*}$ (defined by $D_{e^{\varepsilon^{*}}}\big{(}P_{0}^{B}\parallel P_{1}^{B}\big{)}=\delta$), the function $\delta(\varepsilon)$ decreases rapidly. This rapid decay ensures that even a relatively coarse discretization yields an accurate estimate of $\varepsilon$. Empirical evaluations confirm that $l=O(1)$ is sufficiently accurate in practice (see Section 6), and the overall FFT runtime becomes $O(n\log n)$.
@@
(The GPARVs for BLH, RAPPOR and OUE are asymptotic as $|\mathbb{X}|\to\infty$.) $1-e^{\varepsilon_{0}+\varepsilon}$ $e^{\varepsilon_{0}}-e^{\varepsilon_{0}+\varepsilon}$ $1-e^{\varepsilon}$ 0 $e^{\varepsilon_{0}}-e^{\varepsilon}$ $k$-RR [War65] $\frac{1}{e^{\varepsilon_{0}}+k-1}$ 0 $\frac{k-2}{e^{\varepsilon_{0}}+k-1}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+k-1}$ $\frac{1}{e^{\varepsilon_{0}}+k-1}$ BLH [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+1}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ RAPPOR [EPK14] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\varepsilon_{0}/2}(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $1-e^{-\varepsilon_{0}/2}$ $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ OUE [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2e^{\varepsilon_{0}}(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2}(1-e^{-\varepsilon_{0}})$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ HR [ASZ19] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+1}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
@@
The numerical computation of privacy amplification in the shuffle model has been studied since the introduction of the clone paradigm by Feldman et al. [FMT22]. Prior numerical algorithms can only handle three-point decompositions, such as computing $D_{e^{\varepsilon}}(P_{0}^{C}\parallel P_{1}^{C})$ [FMT22, WPL+24]. However, these techniques do not generalize to decompositions involving more than three points, which are essential for obtaining tight bounds from optimal decompositions.
@@ 5.3 Algorithm for computing privacy amplification lower bounds
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq\delta_{u}.$
@@
$\delta_{l}\leq D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)},$
@@
In previous studies, a common strategy for selecting neighboring datasets $X^{0}$ and $X^{1}$ is to set $x_{2}=x_{3}=\dots=x_{n}$, such that $x_{1}^{0}$, $x_{1}^{1}$, and $x_{2}$ are mutually distinct [FMT23a, WPL+24, GKMP20]. In the special case where $|\mathbb{X}|=2$, the datasets $X^{0}$ and $X^{1}$ can be chosen as
@@
The following theorem provides an efficient method for computing the hockey-stick divergence $D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}$ under this setting. It appeared in [GKMP20].
@@
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
@@
Theorem 14 also provides a method for computing the lower bound in the cases of joint composition and parallel composition. The procedure is analogous to that used for computing the corresponding upper bound: one first computes the decomposition of each sub-local randomizer at inputs $x_{1}^{0},x_{1}^{1},x_{2}$. Since $x_{2}=x_{3}=\cdots=x_{n}$, the resulting decomposition does not involve any left-over distribution. Each component of the decomposition is defined over the set $\{y\mid\mathcal{R}_{i}(x_{1}^{0})(y):\mathcal{R}_{i}(x_{1}^{1})(y):\mathcal{R}_{i}(x_{2})(y)=a:b:c\}$.
@@
While [GKMP20] provided Theorem 14, they explicitly note that their computational method is applicable only when the induced random variable $G$ has support of size three. For example, consider $k$-RR ($k\geq 4$) with datasets $X=(1,3,\dots,3)$ and $X^{\prime}=(2,3,\dots,3)$. The random variable
@@
[…]
@@ 6 Numerical Experiments
We compare our computed bounds for several common local randomizers against existing bounds from prior work. The baselines include two bounds derived from the privacy blanket framework using Hoeffding’s and Bennett’s inequalities, respectively [BBGN19], as well as numerical bounds from the standard clone paradigm [FMT22]. We use the publicly available implementations released by the respective authors.
@@
We present results for five widely used local randomizers: $k$-ary randomized response (with $k=10$[War65], Binary Local Hash (BLH) [WBLJ17], RAPPOR [EPK14], Optimized Unary Encoding (OUE) [WBLJ17], and the Laplace mechanism over $\{0,1\}$ [Dwo06]. For simplicity, we use the asymptotic forms of GPARV for BLH, RAPPOR, and OUE in the regime where the domain size $D\gg 1$. As an illustrative example, we also report how the upper bound for RAPPOR changes with different values of $D$. As shown in Fig. 5, when $D\geq 5$, the upper bound curves become nearly indistinguishable. Furthermore, for $D\geq 3$, where $x_{2}\in\mathbb{X}\setminus\{x_{1}^{0},x_{1}^{1}\}$ can be selected, the lower-bound GPARV for the above local randomizers becomes independent of $D$.

### v6

@@ 5 Efficient Algorithm for Computing the Optimal Bounds
In this section, we present an efficient algorithm to compute the optimal privacy amplification bounds within the general clone framework. Our approach builds upon the previously overlooked concept of the privacy amplification random variable (PARV), originally proposed in the privacy blanket framework [Balle2019]. While PARV offers an exact expression for the amplification bound, the original work did not provide a method for computing it precisely, relying instead on loose analytical approximations.
@@ 5.1 Generalized Privacy amplification random variable
The following theorem is a remarkable result from the privacy blanket paper [Balle2019]. We restate it here in a more general form.
@@
We provide the proof in Appendix C. It is straightforward to observe that Theorem 12 can be directly applied to compute the general clone bound $D_{e^{\varepsilon}}(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}})$. Consider the general clone induced by the primary optimal decomposition; the corresponding random variable $G$ can be expressed as follows:
@@
[…]
$\omega$ and $\gamma$ are defined in Section 3.2.
The quantity $L_{\varepsilon}^{x,x^{\prime}}$ was introduced by Balle et al. [Balle2019] and referred to as the Privacy Amplification Random Variable (PARV). We refer to $G_{\varepsilon}^{x,x^{\prime}}$ as the Generalized Privacy Amplification Random Variable (GPARV). The following theorem follows directly from Theorem 12 and the definition of the primary optimal decomposition in Section 4.2.
@@
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))\leq D_{e^{\varepsilon}}(P_{0}^{B}\|P_{1}^{B})=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
We now briefly revisit how the original privacy blanket paper utilizes the PARV. Using PARV, Balle et al. derived the precise expression for $D_{e^{\varepsilon}}(P_{0}^{B}\|P_{1}^{B})$:
Lemma 3 (Lemma 5.3 in [Balle2019]).
@@
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))\leq D_{e^{\varepsilon}}(P_{0}^{B}\|P_{1}^{B})$ $=\frac{1}{\gamma n}\sum_{m=1}^{n}\binom{n}{m}\gamma^{m}(1-\gamma)^{n-m}\mathbb{E}\left[\sum_{i=1}^{m}L_{i}\right]_{+}$
@@
In the original paper of the privacy blanket framework, a similar property of PARV was provided, but in a loose form [Balle2019]. Specifically, they established that $\gamma(e^{-\varepsilon_{0}}-e^{\varepsilon+\varepsilon_{0}})\leq L\leq\gamma(e^{\varepsilon_{0}}-e^{\varepsilon-\varepsilon_{0}})$. However, the $\varepsilon_{0}$-DP property of $\mathcal{R}$ actually guarantees a tighter bound: $\gamma(1-e^{\varepsilon+\varepsilon_{0}})\leq L\leq\gamma(e^{\varepsilon_{0}}-e^{\varepsilon})$.
@@ 5.2 Algorithm for computing privacy amplification upper bounds
The distributions of $G_{\varepsilon}$ for commonly used local randomizers are summarized in Table 1. Since the optimal decompositions and corresponding GPARVs of BLH, RAPPOR, OUE, and HR depend on the domain size $D$ (see Table 2 in Appendix A), we present their GPARVs in the asymptotic regime (i.e., for $D\gg 1$) for simplicity. The distribution of $G_{\varepsilon}$ for Laplace mechanism on $\{0,1\}$ is presented in Appendix B. (The Laplace mechanism over the interval $[0,1]$ does not satisfy input symmetry.)
@@
Because $G$ is supported on the interval $1-e^{\varepsilon_{0}+\varepsilon},e^{\varepsilon_{0}}-e^{\varepsilon}]$ (Property 2), the discretization step requires $O\left(\frac{e^{\varepsilon_{0}+\varepsilon}}{l}\right)$ operations. For most frequency oracles used on categorical data, $G$ takes values on at most five points (see Table 1), resulting in $O(1)$ discretization time.
@@
The above analysis concerns the error introduced when using the FFT-based algorithm to compute $D_{e^{\varepsilon}}\big(P_{0}^{B},P_{1}^{B}\big)$. From a theoretical perspective, since $\delta$ is typically very small, achieving high-precision computation (i.e., with error $o(\delta)$) would require a very small discretization step size $l$, which incurs a high computational cost.
However, our experimental results show that when discretization is used to infer the smallest value of $\varepsilon$ such that $D_{e^{\varepsilon}}\big(P_{0}^{B},P_{1}^{B}\big)\leq\delta$, the resulting error in $\varepsilon$ is negligible in practice. This is because the distribution of the GPARV is highly sensitive to changes in $\varepsilon$. In particular, once $\varepsilon$ exceeds the true threshold $\varepsilon^{*}$ (defined by $D_{e^{\varepsilon^{*}}}\big(P_{0}^{B},P_{1}^{B}\big)=\delta$), the function $\delta(\varepsilon)$ decreases rapidly. This rapid decay ensures that even a relatively coarse discretization yields an accurate estimate of $\varepsilon$. Empirical evaluations confirm that $l=O(1)$ is sufficiently accurate in practice (see Section 6), and the overall FFT runtime becomes $O(n\log n)$.
@@
(The GPARVs for BLH, RAPPOR, OUE and HR are asymptotic as $|\mathbb{X}|\to\infty$.) $1-e^{\varepsilon_{0}+\varepsilon}$ $e^{\varepsilon_{0}}-e^{\varepsilon_{0}+\varepsilon}$ $1-e^{\varepsilon}$ 0 $e^{\varepsilon_{0}}-e^{\varepsilon}$ $k$-$R$ [11] $\frac{1}{e^{\varepsilon_{0}}+k-1}$ 0 $\frac{k-2}{e^{\varepsilon_{0}}+k-1}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+k-1}$ $\frac{1}{e^{\varepsilon_{0}}+k-1}$ BLH [3] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+1}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ RAPPOR [6] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\varepsilon_{0}/2}(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $1-e^{-\varepsilon_{0}/2}$ $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ OUE [3] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2e^{\varepsilon_{0}}(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2}(1-e^{-\varepsilon_{0}})$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ HR [HR] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+1}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
@@
The numerical computation of privacy amplification in the shuffle model has been studied since the introduction of the clone paradigm by Feldman et al. [Feldman2021]. Prior numerical algorithms can only handle three-point decompositions, such as computing $D_{e^{\varepsilon}}(P_{0}^{C}\|P_{1}^{C})$ [Feldman2021, tv_shuffle]. However, these techniques do not generalize to decompositions involving more than three points, which are essential for obtaining tight bounds from optimal decompositions.
@@ 5.3 Algorithm for computing privacy amplification lower bounds
$D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big)\leq\delta_{u}.$
@@
$\delta_{l}\leq D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big),$
@@
In previous studies, a common strategy for selecting neighboring datasets $X^{0}$ and $X^{1}$ is to set $x_{2}=x_{3}=\dots=x_{n}$, such that $x_{1}^{0}$, $x_{1}^{1}$, and $x_{2}$ are mutually distinct [feldman2023soda, tv_shuffle, Ghazi20]. In the special case where $|\mathbb{X}|=2$, the datasets $X^{0}$ and $X^{1}$ can be chosen as
@@
The following theorem provides an efficient method for computing the hockey-stick divergence $D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big)$ under this setting. It appeared in [Ghazi20].
@@
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
@@
Theorem 14 also provides a method for computing the lower bound in the cases of joint composition and parallel composition. The procedure is analogous to that used for computing the corresponding upper bound: one first computes the decomposition of each sub-local randomizer at inputs $x_{1}^{0},x_{1}^{1},x_{2}$. Since $x_{2}=x_{3}=\cdots=x_{n}$, the resulting decomposition does not involve any left-over distribution. Each component of the decomposition is defined over the set $\{y\mid\mathcal{R}_{i}(x_{1}^{0}):\mathcal{R}_{i}(x_{1}^{1}):\mathcal{R}_{i}(x_{2})=a:b:c\}$.
@@
While [Ghazi20] provided Theorem 14, they explicitly note that their computational method is applicable only when the induced random variable $G$ has support of size three. For example, consider $k$-$R$ ($k\geq 4$) with datasets $X=(1,3,\dots,3)$ and $X^{\prime}=(2,3,\dots,3)$. The random variable
@@
has support over four values $\{e^{\varepsilon_{0}}-e^{\varepsilon},1-e^{\varepsilon_{0}+\varepsilon},1-e^{\varepsilon},\frac{1-e^{\varepsilon}}{e^{\varepsilon_{0}}}\}$. To avoid this, [Ghazi20] assumes inputs of the form $X=(1,C,\dots,C)$ and $X^{\prime}=(2,C,\dots,C)$, with $C$ uniformly distributed over $[k]$. This ensures the above expression has support over only three values: $\{e^{\varepsilon_{0}}-e^{\varepsilon},\ 1-e^{\varepsilon_{0}+\varepsilon},\ 1-e^{\varepsilon}\}$. Our algorithm, by contrast, imposes no such restrictions.
@@ 6 Numerical Experiments
We compare our computed bounds for several common local randomizers against existing bounds from prior work. The baselines include two bounds derived from the privacy blanket framework using Hoeffding’s and Bennett’s inequalities, respectively [Balle2019], as well as numerical bounds from the standard clone paradigm [Feldman2021]. We use the publicly available implementations released by the respective authors.
@@
We present results for five widely used local randomizers: $k$-ary randomized response (with k=10k=10[11], Binary Local Hash (BLH) [3], RAPPOR [6], Optimized Unary Encoding (OUE) [3], and the Laplace mechanism over $\{0,1\}$ [Dwork2006]. For simplicity, we use the asymptotic forms of GPARV for BLH, RAPPOR, and OUE in the regime where the domain size $D\gg 1$. As an illustrative example, we also report how the upper bound for RAPPOR changes with different values of $D$. As shown in Fig. 5, when $D\geq 5$, the upper bound curves become nearly indistinguishable. Furthermore, for $D\geq 3$, where $x_{2}\in\mathbb{X}\setminus\{x_{1}^{0},x_{1}^{1}\}$ can be selected, the lower-bound GPARV for the above local randomizers becomes independent of $D$.
@@
Practical Application.
We validate the improvement brought by our bound on estimation error using the frequency estimation task as an example. Taking 10-RR with $n=1000$ and $\delta=10^{-6}$, we compute, for a target privacy level $\varepsilon$, the value of $\varepsilon_{0}$ required by our bound and by the previously best-known bound (Blanket with Bennett). In the simulation, user data follow a Zipf distribution with parameter $0.7$. Repeating the experiment multiple times, we measure the $\ell_{2}$-error between the estimated and true frequency distributions [^2] For illustration, we adopt the classical unbiased estimation without clipping to $[0,1]$.. The results are summarized in Table 4. As shown, the $\varepsilon_{0}$ suggested by our bound is roughly 10% larger than that given by prior work, leading to a substantial reduction in estimation error.